OpenStudy (anonymous):

Stephen has just built a doghouse.He decides that it needs some carpet to help keep his dog warm during the cold winter months ahead. He measures the rectangular shaped floor of the doghouse and gets a perimeter of 122 inches. If the width is 25 inches what is the area of the carpet that he needs to cover the floor of the doghouse? Please Show Work

OpenStudy (anonymous):

see perimeter of a rectangle = twice the sum of length and width => perimeter of a rectangle = 2(length + width) since you are given perimeter and width, the above equation gives the value of length on solving once you obtain length area = length*width

OpenStudy (anonymous):

2 L + 2 w = 122 2 L + 50 = 122 L = 36 A = 36 * 25 = 900 square inches Ceiling[900/144] = 7 square feet to the next whole square foot. Ceiling[900/(144*9)] = 1 square yard to the next whole square yard